OpenStudy (anonymous):

Help please! y=x^3-3x^2+3x-2 ----Find the extreme values of the function and where they occur.

Directrix (directrix):

By "extreme values," do you want to include relative extrema?

OpenStudy (anonymous):

We could!(: but the question only asks for the extreme values? @directrix

OpenStudy (anonymous):

Math way.com

OpenStudy (anonymous):

it doesn't show me how to do it @Brianawhodat

OpenStudy (anonymous):

Sorry

Directrix (directrix):

y=x^3-3x^2+3x-2 This cubic polynomial enters Quadrant 3. We know that because as x gets smaller and smaller, the dominant term x^3 [think of it as (-x)^3 = -x^3] gets smaller and smaller. The (-, -) quadrant is quad III. This cubic polynomial exits Quadrant I. We know that because as x gets larger and larger, the dominant term x^3 [think of it as (+x)^3 = +x^3] gets larger and larger. The (+,+) quadrant is quad I, If "extreme values" encompasses only end term behavior of the polynomial, then what do you think are the "extreme values?"

OpenStudy (anonymous):

So how do I find the extrema? and where they occur. @Directrix

OpenStudy (anonymous):

I must give points and/or intervals.

Directrix (directrix):

I need to ask @pdd21 a question to answer your question. For the graph shown below, what would you say are the extreme values?